OpenStudy (anonymous):

The measures of two sides of a triangle are 7.3 centimeters and 2.94 centimeters. The triangle's perimeter is 14.8 centimeters. Which equation could not be used to find the length of the third side?

OpenStudy (anonymous):

I assume you have some options, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Please, show.

OpenStudy (anonymous):

7.3 + 2.94 + x = 14.8 14.8 = x + 10.24 14.8 - x = 7.3 + 2.94 x - 14.8 = 2.94 + 7.3

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

yyyyyyyyyyyyyup

OpenStudy (anonymous):

thanks